OpenStudy (anonymous):

what is f(-5) if f(x)=|2x-1|+10? a)21 B)19 C)-1 D-21

jigglypuff314 (jigglypuff314):

plug (-5) in for x into the given equation... so f(-5) = | 2(-5) - 1 | + 10 = | - 10 - 1 | + 10 = | -11 | + 10 absolute values makes the things inside positive = 11 + 10 = answer

OpenStudy (anonymous):

thank you so much!